Show that x = y = z+1 is a secant line of the sphere x² +y² +z²−x−y+z−1 = 0. Also find the intercept made by the sphere on the line.
1
Expert's answer
2020-03-02T17:00:23-0500

The equation of the sphere

"x^2+y^2+z^2-x-y+z-1=0\\
(x-\frac{1}{2})^2+(y-\frac{1}{2})^2+(z+\frac{1}{2})^2=1+\frac{1}{4}+\frac{1}{4}+\frac{1}{4}\\
(x-\frac{1}{2})^2+(y-\frac{1}{2})^2+(z+\frac{1}{2})^2=\frac{7}{4}"

center "O(\frac{1}{2},\frac{1}{2},-\frac{1}{2})" , "R=\sqrt\frac{7}{4}=\frac{\sqrt7}{2}"

Let

"x=y=z+1=t"

now putting the value of "x,y" and "z" in the above equation of the sphere we get

"(t-\frac{1}{2})^2+(t-\frac{1}{2})^2+(t-1+\frac{1}{2})^2=\frac{7}{4}\\
3(t-\frac{1}{2})^2=\frac{7}{4}\\
(t-\frac{1}{2})^2=\frac{7}{12}\\
t-\frac{1}{2}=\sqrt\frac{7}{12} \\and \\t-\frac{1}{2}=-\sqrt\frac{7}{12}\\
t=\frac{1}{2}+\sqrt\frac{7}{12}\\
and\\
t=\frac{1}{2}-\sqrt\frac{7}{12}"

so we get two points where the line cuts the sphere :

"A(\frac{1}{2}+\sqrt\frac{7}{12},\frac{1}{2}+\sqrt\frac{7}{12},-\frac{1}{2}+\sqrt\frac{7}{12})\\
and\\
B(\frac{1}{2}-\sqrt\frac{7}{12},\frac{1}{2}-\sqrt\frac{7}{12},-\frac{1}{2}-\sqrt\frac{7}{12})"

distance between these points is

"AB=\sqrt{(x_B-x_A)^2+(y_B-y_A)^2+(z_B-z_A)^2}"

computable

"(x_B-x_A)^2=(\frac{1}{2}-\sqrt\frac{7}{12}-\frac{1}{2}-\sqrt\frac{7}{12})^2=\\
=(2\sqrt\frac{7}{12})^2=\frac{7}{3}\\
(y_B-y_A)^2=(\frac{1}{2}-\sqrt\frac{7}{12}-\frac{1}{2}-\sqrt\frac{7}{12})^2=\frac{7}{3}\\
(z_B-z_A)^2=(-\frac{1}{2}-\sqrt\frac{7}{12}+\frac{1}{2}-\sqrt\frac{7}{12})^2=\frac{7}{3}"

"AB=\sqrt{\frac{7}{3}+\frac{7}{3}+\frac{7}{3}}=\sqrt{7}=2R"

distance between the two points is equal the diameter of the sphere

hence the given line is a secant the sphere and the intercept made by the line to the sphere is"\sqrt{7}"